OpenStudy (anonymous):

V=150+25(T-80) if v is 500 what is T

OpenStudy (anonymous):

T=94

OpenStudy (sonic):

Basic equation. V=150+25(T-80) -- V = 500 as you defined. 500 = 150 + 25(T-80) 350 = 25(T-80) 14 = T -80 T = 94 -- as anfar says above.